OpenStudy (anonymous):

dy/dx= (x+y+1)^2 Solve the given differential equation by using an appropriate substitution

OpenStudy (anonymous):

\[u=x+y+1\] gives \[u'=1+y'\] Subbing into the equation, you have \[\begin{align*}u'-1&=u^2\\ \frac{du}{u^2-1}&=dx\end{align*}\]

OpenStudy (anonymous):

\(u^2\color{red}+1\), sorry.